OpenStudy (anonymous):

what is the common denominator of 3/9 1/2 and 4/3?

OpenStudy (mathhelp346):

54

OpenStudy (anonymous):

how about 18?

OpenStudy (amistre64):

id add them up and see :)

OpenStudy (amistre64):

\[\frac39+\frac12+\frac43=N\] \[\frac{3(3)}9+\frac{1(3)}2+\frac{4(3)}3=(3)N\] \[1(2)+\frac{1(3)(2)}2+4(2)=3(2)N\] \[2+3+8=6N\] id go with 6

OpenStudy (amistre64):

for lowest common denom :)